# GIVT: Generative Infinite-Vocabulary Transformers

### v2

@@ Acknowledgments
We would like to thank André Susano Pinto, Neil Houlsby, Eirikur Agustsson, and Lucas Theis for inspiring discussions and helpful feedback on this project. We also thank Han Zhang for support with the VAE training code.
Appendix A Training details
A.1 Loss function and implementation details
@@
We start with the case where GIVT models the distribution of each feature vector or soft token in the sequence independently across channels with a scalar $k$-mixture GMM, conditionally on all previous feature vectors in the sequence. The predicted GIVT output for a mini batch of size $B$ of $d$-dimensional feature sequences of length $L$ is a $B\times L\times 3dk$ tensor $y=[m;s;\pi]$ where $m,s$ and $\pi$ are $B\times L\times dk$ tensors stacked along the last dimension comprising means, standard deviations, and mixture weights, respectively. We assume the entries of $s$ to be lower-bounded by a small positive constant $\epsilon=10^{-5}$ (e.g., by applying a softplus function to the corresponding network outputs and clipping to $\epsilon$) and the entries of $\pi$ to be non-negative. $m$ is composed as $[m^{(1)};\ldots;m^{(k)}]$ where the $m^{(n)}$ are the $B\times L\times d$ mean tensors of the GMM components stacked along the last axis; $s$ and $\pi$ are composed in the same way. The mixture weights are assumed normalized across components (e.g., via a softmax function), i.e., $\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}=1$ for all $b,\ell,c$.
@@
Then, the negative log-likelihood of the feature sequences $\{z_{b}\}_{b=1}^{B}$ (i.e. the training images after embedding via the VAE encoder and applying reparametrization) w.r.t. the distribution $\tilde{p}$ predicted by GIVT can be written as
@@
$=-\sum_{b=1}^{B}\log\left(\prod_{\ell=1}^{L}\prod_{c=1}^{d}\left(\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right)\right)$
$=-\sum_{b=1}^{B}\sum_{\ell=1}^{L}\sum_{c=1}^{d}\log\left(\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right),$ (2)
@@
When using a single Gaussian instead of a GMM to model each output channel (i.e. $k=1$, as in most experiments) Eq. (2) becomes
@@
A.2 Image generation
For the image generation experiments on ImageNet, we adapt the CNN-based VQ-GAN tokenizer from MaskGIT (see vqgan_tokenizer.py). We replace the VQ layer and related losses with a Gaussian reparametrization layer [27], and we use the hyper parameters given in Table 5 for the CNN. For our transformers, we use the standard pre-LN setup (see, e.g., UViM vtt.py for a concrete implementation). We use the BERT-Large [12] configuration for the main models (24 layers, 16 attention heads, MLPs with hidden dimension 4096, embedding dimension 1024).
@@ A.2 Image generation
A.3 Panoptic Segmentation and Depth Estimation
@@ Appendix B DB-CFG implementation
Appendix C Unconditional image generation
@@
[…]
ADM [13] $\numprint{26.2}$0
[…]
GIVT-Causal (Ours) $\numprint{24.95}$
[…]
Table 6: Results for unconditional $256\times 256$ ImageNet generation. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. We obtained samples for MAGE [36] using their GitHub code. Note that the MAGE Base and Large models have considerably more model parameters than our corresponding models because MAGE has a latent encoder-decoder model (rather than a decoder-only model).
Appendix D Probing intermediate representations
We extract intermediate representations form a GIVT-causal trained for $256\times 256$ unconditional ImageNet generation by average-pooling the output of a given layer over the sequence dimension (which results in a feature vector of dimension 1024 for L-sized models). Thereby, the input image is first encoded with the VAE encoder and then fed to GIVT as during teacher forcing (i.e. the latent sequence is first right-shifted and padded, and then fed to GIVT with causal attention mask).
@@ Appendix D Probing intermediate representations
We choose layer 18 for the linear probing experiments using the full ImageNet training set. To select hyperparameters we set aside 1% of the training data and consider the selection of schedules and hyperparameters described in [62, App. A]. The results, along with other generative models from the literature, can be found in Table 3. See Sec. 5.1 for a discussion of these results.
Appendix E Additional visual examples
In Fig. 10, we show reconstructions from our VAE when feeding it images from the ImageNet validation set. We see low-level texture variations in the reconstructions when sampling from the encoder distribution multiple times. Fig. 11 shows multiple samples for a fixed label to show the diversity of our model and compare with baselines. In Fig. 12 we show various samples from MaskGIT for the same classes that we shown in Fig. 2. Fig. 13 shows samples from our $512\times 512$ GIVT-MaskGIT model. We explore changing labels midway through sampling (i.e., after generating the top half of the image) for GIVT-Causal in Fig. 14.
For our GIVT-Casual UViM models and VQ baselines, we show visual outputs in Figs. 15 16.
@@ Appendix E Additional visual examples
  points = p_cond.loc[None, ...] + jnp.linspace(-10, 10, 1001).reshape(
      -1, 1, 1, 1)
@@
Figure 9: jax.jit-compatible implementation of the rejection sampler used for DB-CFG
@@
GIVT-Causal (Ours)
GIVT-MaskGIT (Ours)
Validation Set
BigGAN [5] (via [6, Fig. 9])
VQ-based MaskGIT [6, Fig. 9]
Figure 11: Various samples for the 980 class to show the diversity of our samples. For reference, we copy examples from from [6, Fig. 9].
Figure 12: Selected $256\times 256$ samples from GIVT-MaskGIT, for the same 10 ImageNet classes as in Fig. 2.
Figure 13: Selected $512\times 512$ samples from the GIVT-MaskGIT, for the same 10 ImageNet classes as in Fig. 2.
@@
Input
Ground Truth
GIVT-Causal UViM
VQ-based UViM
@@
Input
Ground Truth
GIVT-Causal UViM
VQ-based UViM

### v4

@@ Acknowledgments
We would like to thank André Susano Pinto, Neil Houlsby, Eirikur Agustsson, Lucas Theis, and Basil Mustafa for inspiring discussions and helpful feedback on this project. We also thank Han Zhang for support with the VAE training code.
Appendix 0.A Additional discussions
Why continuous latents for images?  Continuous latents naturally fit intrinsically continuous-valued data such as image representations, and avoid excessive compression and the resulting information loss induced by VQ. Removing VQ also avoids the well-documented challenges of stochastic-gradient-based discrete representation learning: VQ-VAE [49] and its variations require a VQ optimization problem, which on its own is NP-hard, to be solved jointly with an embedding learning problem. A large body of literature [7, 27, 33, 37, 49] focuses on mitigating issues resulting from these optimization difficulties, such as low vocabulary utilization (Sec. 1). For example, [46] avoids vector quantization by using a product codebook, but still relies on non-differentiable scalar quantization. In contrast, with GIVT we obtain SOTA results based on a $\beta$-VAE which does not include any non-differentiable operations and does not require any advanced tricks from the VQ literature [7, 27, 33, 37, 49].
Continuous latents vs. infinite vocabulary  In theory, continuous (i.e., real-valued) latents/tokens always imply an infinite vocabulary when mapping them exactly to a discrete code. In other words, the infinite vocabulary is a consequence of continuous latents. In practice, continuous latents are represented with finite precision, e.g., float32, but using the raw 32-bit representation as a discrete code would still imply an impractically large vocabulary size $2^{32}\approx 4.3$B for a single latent dimension. GIVT directly models the continuous latents and avoids materializing this vocabulary.
Appendix 0.B Architecture details
Architecture details for different model variants used for image generation are listed in Table 5. We rely on the standard pre-LN setup (see, e.g., UViM vtt.py for a concrete implementation). For the UViM experiments, we use GIVT-Causal with the Default config and a cross-attention layer inserted after every self-attention layers as in [33] to fuse visual features extracted from the input RGB image.
Adapters are constructed by stacking 8 convolutional, bijective iRevNet blocks [28] with hidden channel dimension $4d$ (resulting in 112k additional parameters for $d=16$). Each block consists of 3 Conv layers, interleaved with GroupNorm and ReLU, and has identical input and output shapes $w\times h\times d$ (i.e., the adapter is applied to the VAE latent $z$ before reshaping it into a sequence). We base our iRevNet block on the reference implementation iRevNet.py, replacing BatchNorm with GroupNorm and removing subsampling layers.
Table 5: Architecture details for different model variants for image generation. We also specify the KL weight $\beta$ used to train the associated VAE. The Base architecture is used to explore the feature dimension $d$, number of mixtures $k$ and $\beta$ (Fig. 5). We use $\beta=5\cdot 10^{-5}$ by default, and $\beta=10^{-5}$ for Large models. GIVT-Causal-L at 512px relies on a space-to-depth transformation, where pairs of consecutive 16-dimensional feature vectors are stacked into 32-dimensional vectors, resulting in sequence length of 512 instead of 1024.
Model Size Res. $d$ $k$ Width Depth MLP Heads Param. Tok. Drop. $\beta$
Causal Base 256 $4\ldots 32$ $1\ldots 32$ 768 12 3072 12 86M 256 0.1 $0.25\ldots 2\cdot 10^{-4}$
Causal Default 256 16 16 1024 24 4096 16 304M 256 0.2 $5\cdot 10^{-5}$
Causal Large 256 16 16 1536 48 8192 16 1.67B 256 0.3 $10^{-5}$
Causal Large 512 32 32 1536 48 8192 16 1.67B 512 0.1 $10^{-5}$
MaskGIT Default 256 16 16 1024 24 4096 16 304M 256 0.4 $5\cdot 10^{-5}$
MaskGIT Default 512 16 16 1024 24 4096 16 304M 1024 0.4 $5\cdot 10^{-5}$
Figure 8: Loss (NLL) curves when training the Base-sized GIVT-causal models from Fig. 5, with latent dimension $d=16$, as a function of the number of mixtures $k$ and the KL weight $\beta$ in the $\beta$-VAE. Reducing $\beta$ or increasing $k$ leads to a reduction in loss, which not always translates into a lower FID. The loss is computed using distrax as described in Sec. 0.C.1. Alternative implementations can lead to scaling by $\frac{1}{d}$ ($\frac{1}{16}$ here).
Appendix 0.C Training details
0.C.1 Loss function and implementation details
@@
We start with the case where GIVT models the distribution of each feature vector or soft token in the sequence, conditionally on all previous feature vectors in the sequence, with a $k$-mixture GMM whose components are independent across channels (i.e., the components are multivariate Gaussian distributions with diagonal covariance matrix). The predicted GIVT output for a mini batch of size $B$ of $d$-dimensional feature sequences of length $L$ is a $B\times L\times(2dk+k)$ tensor $y=[m;s;\pi]$ where $m,s$ are $B\times L\times dk$ tensors and $\pi$ is a $B\times L\times k$ tensor, all stacked along the last dimension, comprising means, standard deviations, and mixture weights, respectively. We assume the entries of $s$ to be lower-bounded by a small positive constant $\epsilon=10^{-5}$ (e.g., by applying a softplus function to the corresponding network outputs and clipping to $\epsilon$) and the entries of $\pi$ to be non-negative. $m$ is composed as $[m^{(1)};\ldots;m^{(k)}]$ where the $m^{(n)}$ are the $B\times L\times d$ mean tensors of the GMM components stacked along the last axis; $s$ is composed in the same way. The mixture weights are assumed normalized across components (e.g., via a softmax function), i.e., $\sum_{n=1}^{k}\pi_{b,\ell}^{(n)}=1$ for all $b,\ell$.
@@
Then, the negative log-likelihood (NLL) of the feature sequences $\{z_{b}\}_{b=1}^{B}$ (i.e. the training images after embedding via the VAE encoder and applying reparametrization) w.r.t. the distribution $\tilde{p}$ predicted by GIVT can be written as
@@
$=-\sum_{b=1}^{B}\log\left(\prod_{\ell=1}^{L}\left(\sum_{n=1}^{k}\pi_{b,\ell}^{(n)}\prod_{c=1}^{d}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right)\right)$
$=-\sum_{b=1}^{B}\sum_{\ell=1}^{L}\log\left(\sum_{n=1}^{k}\pi_{b,\ell}^{(n)}\prod_{c=1}^{d}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right),$ (2)
@@
When using a single Gaussian instead of a GMM to model each output channel (i.e. $k=1$ s.t. $\pi_{b,\ell}^{(1)}=1$ for all $b,\ell$) Eq. (2) becomes
@@
Fig. 8 shows the training loss curves as a function of VAEs with different KL weights $\beta$ and $k$, following the setup of Fig. 5. Increasing $k$ and reducing $\beta$ leads to lower loss values. Note, however, that a lower loss does not always lead to a lower sampling FID (see Fig. 5). Importantly, initial and final loss values, as well as the relative reduction in training loss throughout training can differ significantly from the values typically observed for the discrete cross-entropy or NLL, e.g., in language modeling.
@@
Finally, loss computation and sampling can easily be implemented with dedicated deep learning packages, for example the JAX library distrax [12]:
# With z and m, s, pi as in Eq. (2)
pdf = distrax.MixtureSameFamily(
mixture_distribution=distrax.Categorical(logits=pi),
components_distribution=distrax.MultivariateNormalDiag(
loc=m, scale_diag=s))
# Sample from next token distribution (teacher forcing)
samples = pdf.sample()
# Compute NLL
loss = -pdf.log_prob(z).mean()
0.C.2 Image generation
For the image generation experiments on ImageNet, we adapt the CNN-based VQ-GAN tokenizer from MaskGIT (see vqgan_tokenizer.py). We replace the VQ layer and related losses with a Gaussian reparametrization layer [30], and we use the hyper parameters given in Table 6 for the CNN. See Sec. 0.B for details on the GIVT model architecture and variants.
@@ 0.C.2 Image generation
Table 7: Results for unconditional $256\times 256$ ImageNet generation. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. We obtained samples for MAGE [39] using their GitHub code. The MAGE Large models have considerably more model parameters than our GIVT-Causal models because MAGE has a latent encoder-decoder model (rather than a decoder-only model).
[…]
ADM [14] $\numprint{26.2}$0 $\numprint{0.61}$ $\numprint{0.63}$
[…]
@@
0.C.3 Panoptic Segmentation and Depth Estimation
@@ Appendix 0.D DB-CFG implementation
Appendix 0.E Unconditional image generation
@@
Figure 10: Effect of different sampling strategies and model variants (GIVT-Causal-L) on precision and recall, complementing Fig. 5 for FID. Increasing the number of mixtures $k$ and adding an adapter have lead to compounding effects. DB-CFG is the most effective sampling strategy to increase precision for all model variants.
Appendix 0.F Probing intermediate representations
We extract intermediate representations form a GIVT-causal trained for $256\times 256$ unconditional ImageNet generation by average-pooling the output of a given layer over the sequence dimension (which results in a feature vector of dimension 1024). Thereby, the input image is first encoded with the VAE encoder and then fed to GIVT as during teacher forcing (i.e. the latent sequence is first right-shifted and padded, and then fed to GIVT with causal attention mask).
@@ Appendix 0.F Probing intermediate representations
We choose layer 18 for the linear probing experiments using the full ImageNet training set. To select hyperparameters we set aside 1% of the training data and consider the selection of schedules and hyperparameters described in [68, App. A]. The results, along with other generative models from the literature, can be found in Table 4. See Sec. 5.1 for a discussion of these results.
Appendix 0.G Comparison of methods in terms of FLOPs
Among the models reported in Table 1 only the two diffusion models [14, 51] report FLOPs, but not the sequence or masked models, which are most closely related to our GIVT models. Assuming the same autoregressive transformer optimizations (e.g. caching) and VAE inference cost, we obtain the following approximate comparison, using the $N$ FLOPs of GIVT-Causal-L+A as a reference. MaskGIT and GIVT-MaskGIT require virtually the same number of FLOPs for inference.
Table 8: Approximate comparison in FLOPs of GIVT and autoregressive baselines from the literature.
Model Guidance Steps Params. FLOPs FID
GIVT-Causal-L+A 256 1.67B $N$ 3.46
VQGAN [20] 256 1.4B $\simeq N$ 17.04
ViT-VQGAN-L [76] 1024 1.7B $\simeq 4N$ 4.17
GIVT-Causal-L+A DB-CFG 256 1.67B $\simeq 2N$ 2.59
VQGAN [20] CG $=0.05$ 256 1.4B $\simeq 20N$ 5.20
ViT-VQGAN-L [76] CG $=0.5$ 1024 1.7B $\simeq 8N$ 3.04
Appendix 0.H Broader impact
The method described in this paper modifies transformer decoder-only models to enable generation of sequences of feature vectors, which can be seen as a research endeavor on the foundation of transformer models, with many potential downstream applications. The broader impact will depend largely on the downstream application, and how it is affected by this work.
Here, we focus on generation of visual data. For class-conditional image generation, we rely on ImageNet, and the resulting models provide basic control of the generated image content via the class label. These models do not allow for fine-grained control of image content, or manipulation of existing images. The biases and issues inherent with ImageNet are well-studied and documented [Uday Prabhu and Birhane, 2020]. We expect that our image generation models could reflect some of these issues, and we advise users to use and deploy the model carefully, taking these issues into consideration.
It is relatively straight-forward to extend our image generation models to a text-to-image interface, which enables training on much larger and less curated image/text data sets collected from the web. Such models allow for much more fine-grained control and hence could be misused for malicious purposes such as spreading misinformation or identity theft. Furthermore, such models might reflect biases in large non-curated data sets from the web, such as harmful stereotypes. The deployment and release of such models thus has to be handled with even more care.
@@ Appendix 0.H Broader impact
points = p_cond.loc[None, …] + jnp.linspace(-10, 10, 1001).reshape(
-1, 1, 1, 1)
@@
Figure 11: jax.jit-compatible implementation of the rejection sampler for DB-CFG.
@@
Appendix 0.I Additional visual examples
In Fig. 12, we show reconstructions from our VAE when feeding it images from the ImageNet validation set. We see low-level texture variations in the reconstructions when sampling from the encoder distribution multiple times. Fig. 13 shows multiple samples for a fixed label to show the diversity of our model and compare with baselines. In Fig. 14 and 16 we present samples from different $256\times 256$ GIVT-Causal variants. In Fig. 17 we show samples from MaskGIT for the same classes that we shown in Fig. 1. Fig. 18 shows samples from our $512\times 512$ GIVT-MaskGIT model. We explore changing labels midway through sampling (i.e., after generating the top half of the image) for GIVT-Causal in Fig. 19.
For our GIVT-Casual UViM models and VQ baselines, we show visual outputs in Figs. 21 and 21.
GIVT-Causal (Ours) GIVT-MaskGIT (Ours) Validation Set
BigGAN [5] (via [7, Fig. 9]) MaskGIT (VQ) [7, Fig. 9]
Figure 13: Various samples for the 980 class to show the diversity of our samples (without DB-CFG). For reference, we copy examples from from [7, Fig. 9].
Figure 14: Selected $256\times 256$ samples from the GIVT-Causal ($t=0.95$, DB-CFG $=0.4$), for the same 10 ImageNet classes as in Fig. 1.
VAE Sample Step 1 Step 4 Step 8 Step 16
Figure 16: Selected $256\times 256$ samples from GIVT-Causal-L+A ($t=0.95$, DB-CFG $=0.4$), for the same 10 ImageNet classes as in Fig. 1.
Figure 17: Selected $256\times 256$ samples from GIVT-MaskGIT ($t_{C}=60$, DB-CFG $=0.1$), for the same 10 ImageNet classes as in Fig. 1.
Figure 18: Selected $512\times 512$ samples from the GIVT-MaskGIT ($t_{C}=140$), for the same 10 ImageNet classes as in Fig. 1.
@@ Appendix 0.I Additional visual examples
Input Ground Truth GIVT-Causal UViM VQ-based UViM
@@
Input Ground Truth GIVT-Causal UViM VQ-based UViM